OpenStudy (anonymous):

in a recent year the total sales at restaurant A were 325,000 more than restaurant B. Total sale at the two stores were 600,000,000. What were the sales for the year at each store?

OpenStudy (anonymous):

A = B + 325,000 A + B = 600,000,000 so B + 325,000 + B = 600,000,000 2B + 325,000 = 600,000,000 Solve for B, then plug B into the original A equation to solve for A.

OpenStudy (anonymous):

ok i got it thank you